# Global Existence for General Systems of Isentropic Gas Dynamics via a Weighted Pressure Perturbation Approach

### v1

@@ Abstract
This paper establishes the global existence of bounded entropy solutions to the one-dimensional system of isentropic gas dynamics with general pressure laws, allowing for the presence of vacuum states. We introduce a novel regularization scheme based on a weighted pressure perturbation of the form $\tilde{P}(\rho,\delta)=\int_{\delta}^{\rho}s^{-2}(s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta))\,ds$. Unlike previous methods that modify the continuity equation, our approach strictly preserves mass conservation while creating a degenerate hyperbolic boundary at the cut-off density $\rho=\delta$. We provide a rigorous derivation of the eigenstructure and Riemann invariants for the perturbed system. A key theoretical contribution is the proof that the perturbed pressure exactly inherits the convexity properties of the original state equation ($2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}=2P^{\prime}+\rho P^{\prime\prime}$), thereby ensuring the strict convexity of the mechanical energy entropy without additional assumptions. Using the method of vanishing viscosity, we derive uniform $L^{\infty}$ bounds via invariant regions and establish the $H^{-1}_{\mathrm{loc}}$ compactness of entropy dissipation measures. The global existence for the vacuum-free limit is then proved using the theory of compensated compactness and singular perturbation analysis of the entropy equations.
@@ 1 Introduction
The main challenge in analyzing (1.1) arises from the degeneracy of hyperbolicity at the vacuum state $\rho=0$. To overcome this difficulty while strictly preserving the mass conservation law, we propose to approximate (1.1) by the following perturbed system with a weighted pressure term:
@@
While inspired by Lu’s pioneering approach [4], we propose an alternative regularization technique that modifies only the pressure term, thereby strictly preserving the exact mass conservation law $\rho_{t}+(\rho u)_{x}=0$. For a parameter $\delta>0$, we introduce a novel weighted pressure perturbation:
$\tilde{P}(\rho,\delta)=\int_{\delta}^{\rho}\left(\frac{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}{s^{2}}\right)ds,\quad\rho\geq\delta.$ (1.8)
This formulation ensures that the acoustic characteristics vanish exactly at $\rho=\delta$ (i.e., $\tilde{P}^{\prime}(\delta)=0$), creating a sequence of strictly hyperbolic systems that approximate (1.1). Our approach maintains the geometric structure of the invariant regions required for the $L^{\infty}$ estimates, but does so within a framework that respects the original continuity equation. The detailed motivation and construction of the perturbation function are provided in Appendix A, while a comparative analysis with Lu [4] is presented in Appendix B.
@@ 2.1 Properties of the Weighted Pressure
Defining the squared sound speed for the perturbed system as $\tilde{c}^{2}(\rho):=\partial_{\rho}\tilde{P}(\rho,\delta)$, we obtain from (1.3):
$\tilde{c}^{2}(\rho)=P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}=\frac{\rho^{2}P^{\prime}(\rho)-\delta^{2}P^{\prime}(\delta)}{\rho^{2}}.$ (2.2)
@@
1. Let $g(\rho)=\rho^{2}P^{\prime}(\rho)$. Compute $g^{\prime}(\rho)=2\rho P^{\prime}(\rho)+\rho^{2}P^{\prime\prime}(\rho)=\rho(2P^{\prime}+\rho P^{\prime\prime})>0$. Since $\rho>\delta$, $g(\rho)>g(\delta)$, which implies $\rho^{2}P^{\prime}(\rho)>\delta^{2}P^{\prime}(\delta)$. Thus, $\tilde{c}^{2}(\rho)>0$.
2. We compute the convexity term for the perturbed pressure. Note that $\tilde{P}^{\prime}=\tilde{c}^{2}$ and $\tilde{P}^{\prime\prime}=(\tilde{c}^{2})^{\prime}$. We calculate:
@@
$=(2P^{\prime}(\rho)+\rho P^{\prime\prime}(\rho))-\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}+\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}$
@@
Remarkably, the singular terms cancel exactly. Since the original pressure satisfies $2P^{\prime}+\rho P^{\prime\prime}>0$, the perturbed pressure satisfies the exact same condition. ∎
@@ 2.2 Eigenstructure and Riemann Invariants
The system (1.1) with pressure $\tilde{P}$ can be written in quasilinear form $\mathbf{U}_{t}+A(\mathbf{U})\mathbf{U}_{x}=0$. The eigenvalues are:
@@
The Riemann invariants $z,w$ correspond to quantities constant along integral curves of the eigenvectors $\mathbf{r}_{1},\mathbf{r}_{2}$. They satisfy $du=\pm\frac{\tilde{c}(\rho)}{\rho}d\rho$. Defining
$z(\rho,u)=u-\int_{\delta}^{\rho}\frac{\sqrt{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}}{s^{2}}\,ds,\quad w(\rho,u)=u+\int_{\delta}^{\rho}\frac{\sqrt{s^{2}P^{\prime}(s)-\delta^{2}P^{\prime}(\delta)}}{s^{2}}\,ds.$ (2.3)
The integral is well-defined near $\rho=\delta$ from (2.2) since $\tilde{c}^{2}(\delta)=0$ and $(\tilde{c}^{2})^{\prime}(\delta)>0$.
Remark 2.2 (Consistency with Original Riemann Invariants).
The Riemann invariants $z,w$ of the perturbed system are constructed to be consistent with the classic Riemann invariants $R$ of the original isentropic Euler equations (1.1) in the limit $\delta\to 0$.
The classic Riemann invariants $R$ are defined by $R(\rho)=\int_{0}^{\rho}\frac{c(s)}{s}\,ds$, where $c(s)=\sqrt{P^{\prime}(s)}$. We show that $\lim_{\delta\to 0}\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds=\int_{0}^{\rho}\frac{c(s)}{s}\,ds$.
The integral is given by:
$I(\rho,\delta)=\int_{\delta}^{\rho}\frac{\sqrt{P^{\prime}(s)-\frac{\delta^{2}P^{\prime}(\delta)}{s^{2}}}}{s}\,ds.$
Since $P$ satisfies the vacuum singularity condition (A3), the classic integral $\int_{0}^{\rho}\frac{c(s)}{s}\,ds$ converges. For any fixed $\rho_{1}\in(0,\rho)$, we decompose the integral:
$I(\rho,\delta)=\int_{\delta}^{\rho_{1}}\frac{\tilde{c}(s)}{s}\,ds+\int_{\rho_{1}}^{\rho}\frac{\tilde{c}(s)}{s}\,ds.$
  1. 1.
    Interval $\rho_{1},\rho]$: For $\in[\rho_{1},\rho], ss$ is bounded away from 0. Thus, $\lim_{\delta\to 0}\frac{\delta^{2}P^{\prime}(\delta)}{s^{2}}=0$, and $\tilde{c}(s)$ converges uniformly to $c(s)=\sqrt{P^{\prime}(s)}$. By the Lebesgue Dominated Convergence Theorem:
    $\lim_{\delta\to 0}\int_{\rho_{1}}^{\rho}\frac{\tilde{c}(s)}{s}\,ds=\int_{\rho_{1}}^{\rho}\frac{c(s)}{s}\,ds.$
  2. 2.
    Interval $\delta,\rho_{1}]$: For $s\in[\delta,\rho_{1}]$, since $\tilde{c}(\delta)=0$ and the integrand $\frac{\tilde{c}(s)}{s}$ is always non-negative and bounded above by $\frac{c(s)}{s}$, the limit of this portion converges to the remainder of the classic integral:
    $\lim_{\delta\to 0}\int_{\delta}^{\rho_{1}}\frac{\tilde{c}(s)}{s}\,ds=\int_{0}^{\rho_{1}}\frac{c(s)}{s}\,ds.$
Combining the limits, we conclude:
$\lim_{\delta\to 0}I(\rho,\delta)=\int_{0}^{\rho_{1}}\frac{c(s)}{s}\,ds+\int_{\rho_{1}}^{\rho}\frac{c(s)}{s}\,ds=\int_{0}^{\rho}\frac{c(s)}{s}\,ds.$
This confirms that the regularization scheme correctly recovers the physical Riemann invariants in the non-degenerate limit.
Proposition 2.3 (Explicit Riemann Invariants for Perturbed Polytropic Gas).
For the perturbed system with pressure $\tilde{P}$ and $P(\rho)=\kappa\rho^{\gamma}$, the Riemann invariants $z,w$ are defined by:
$z,w=u\mp\sqrt{\kappa\gamma}\int_{\delta}^{\rho}\frac{\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}}{s^{2}}\,ds.$ (2.4)
@@
The Riemann invariants $z$ and $w$ are defined by integrating the characteristic form of the equations:
$z,w=u\mp\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds,$
where $\tilde{c}(s)=\sqrt{\tilde{P}^{\prime}(s)}$ is the perturbed sound speed. For the polytropic gas $P(\rho)=\kappa\rho^{\gamma}$, we have $P^{\prime}(\rho)=\kappa\gamma\rho^{\gamma-1}$. Using (2.2):
$\tilde{P}^{\prime}(s)$ $=P^{\prime}(s)-\frac{\delta^{2}P^{\prime}(\delta)}{s^{2}}$
$=\kappa\gamma s^{\gamma-1}-\kappa\gamma\frac{\delta^{\gamma+1}}{s^{2}}$
$=\frac{\kappa\gamma}{s^{2}}(s^{\gamma+1}-\delta^{\gamma+1}).$
The perturbed sound speed is:
$\tilde{c}(s)=\sqrt{\tilde{P}^{\prime}(s)}=\frac{\sqrt{\kappa\gamma}}{s}\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}.$
Substituting this into the integral yields the result. ∎
Remark 2.4 (On the Conservation of Mass and Eigenvalue Degeneracy).
It is worth discussing why the regularization scheme in [4] requires a modification of the continuity equation, whereas our approach does not.
In [4], the continuity equation is perturbed as $\rho_{t}+(\rho u-2\delta u)_{x}=0$. The inclusion of the term $2\delta u$ essentially shifts the convective flux. This specific modification was designed to match the structure of the pressure perturbation $P_{1}(\rho)\sim\int(1-2\delta/\rho)P^{\prime}d\rho$. The algebraic goal in [4] was to force the acoustic component of the eigenvalues to factorize as:
$\lambda_{1,2}^{Lu}=u\mp\frac{\rho-2\delta}{\rho}\sqrt{P^{\prime}(\rho)}.$
The factor $\frac{\rho-2\delta}{\rho}$ ensures that the strictly hyperbolic character degenerates exactly at the boundary $\rho=2\delta$, creating an invariant region. Without modifying the mass flux, Lu’s specific pressure construction would not yield such a clean degeneracy in the eigenvalues.
In contrast, our approach recognizes that the degeneracy of hyperbolicity depends solely on the vanishing of the sound speed. For the system with exact mass conservation $\rho_{t}+(\rho u)_{x}=0$, the eigenvalues are simply $\lambda_{1,2}=u\mp\sqrt{\tilde{P}^{\prime}(\rho)}$. Therefore, to achieve the requisite boundary degeneracy, it suffices to construct a pressure $\tilde{P}$ such that $\tilde{P}^{\prime}(\delta)=0$. Our weighted pressure perturbation (1.3) satisfies this condition directly:
$\tilde{P}^{\prime}(\delta)=P^{\prime}(\delta)-\frac{\delta^{2}P^{\prime}(\delta)}{\delta^{2}}=0.$
Consequently, $\lambda_{1,2}\to u$ as $\rho\to\delta$, creating the necessary ”hard” boundary for the invariant regions without the need to violate the physical law of mass conservation.
@@ 3 Entropy Analysis
To apply compensated compactness, we require a strictly convex entropy pair $(\eta,q)$ that satisfies the entropy identity $\partial_{t}\eta(\mathbf{U})+\partial_{x}q(\mathbf{U})=0$ for smooth solutions.
@@ 3.1 Construction of the Mechanical Energy Entropy Pair
The mechanical energy is the canonical entropy for the isentropic system. We define the specific internal energy $e(\rho)$ for the perturbed pressure via the thermodynamic relation $de=-\tilde{P}d(1/\rho)=\frac{\tilde{P}}{\rho^{2}}d\rho$. Integrating from the cut-off density, we obtain:
@@
The mechanical energy entropy (per unit volume) is $\eta^{*}=\frac{1}{2}\rho u^{2}+\rho e(\rho)$. In terms of the conservative variables $(\rho,m)$ with $m=\rho u$, this is:
@@
$q^{*}(\rho,m)=\left(\eta^{*}+\tilde{P}(\rho)\right)u.$ (3.3)
@@
We compute the derivatives of $\eta^{*}$ with respect to the conservative variables. First, note that from the definition of $e(\rho)$, we have $e^{\prime}(\rho)=\frac{\tilde{P}(\rho)}{\rho^{2}}$. Thus,
@@
The compatibility condition for an entropy pair is $\nabla_{U}\eta^{*}\cdot F^{\prime}(U)=\nabla_{U}q^{*}$. Alternatively, for smooth solutions:
$\partial_{t}\eta^{*}+\partial_{x}q^{*}$ $=\eta^{*}_{\rho}(\rho_{t}+(\rho u)_{x})+\eta^{*}_{m}((\rho u)_{t}+(\rho u^{2}+\tilde{P})_{x})$
$=\eta^{*}_{\rho}(0)+\eta^{*}_{m}(0)=0.$
Thus, $(\eta^{*},q^{*})$ is a valid entropy-entropy flux pair. ∎
@@
The entropy $\eta^{*}(\rho,m)$ is strictly convex for $\rho>\delta$.
@@
The Hessian matrix $\nabla^{2}\eta^{*}$ is given by:
@@
We compute the second derivatives:
$\eta^{*}_{mm}=\frac{1}{\rho},\quad\eta^{*}_{\rho m}=-\frac{m}{\rho^{2}}=-\frac{u}{\rho}.$
For $\eta^{*}_{\rho\rho}$, we differentiate $\eta^{*}_{\rho}$:
$\eta^{*}_{\rho\rho}$ $=\frac{\partial}{\partial\rho}\left(-\frac{m^{2}}{2\rho^{2}}+e(\rho)+\frac{\tilde{P}}{\rho}\right)$
$=\frac{m^{2}}{\rho^{3}}+e^{\prime}(\rho)+\frac{\tilde{P}^{\prime}\rho-\tilde{P}}{\rho^{2}}$
$=\frac{u^{2}}{\rho}+\frac{\tilde{P}}{\rho^{2}}+\frac{\tilde{P}^{\prime}}{\rho}-\frac{\tilde{P}}{\rho^{2}}$
@@
$=\frac{u^{2}}{\rho^{2}}+\frac{\tilde{P}^{\prime}}{\rho^{2}}-\frac{u^{2}}{\rho^{2}}$
@@
[…]
@@
This is due to the constructed degeneracy of the acoustic speed at the cut-off density: $\tilde{P}^{\prime}(\delta)=0$. Consequently, the determinant of the Hessian vanishes at the boundary:
@@
The entropy $\eta^{*}$ is therefore a degenerate convex function at the boundary $\rho=\delta$, consistent with the transition to parabolic degeneracy.

### v2

@@ Abstract
This paper establishes the global existence of bounded entropy solutions to the one-dimensional system of isentropic gas dynamics with general pressure laws, allowing for the presence of vacuum states. We introduce a novel regularization scheme based on a weighted pressure perturbation designed to enforce acoustic degeneracy at a cut-off density $\rho=\delta$. A distinguishing feature of our approach is the strict preservation of the mass conservation law, in contrast to previous regularization methods that require modifying the continuity equation. We prove that the perturbed pressure $\tilde{P}$ exactly inherits the convexity structure of the original equation of state (satisfying $2\tilde{P}^{\prime}+\rho\tilde{P}^{\prime\prime}=2P^{\prime}+\rho P^{\prime\prime}$), thereby ensuring the strict convexity of the mechanical entropy without auxiliary transformations. This structural preservation allows for the direct application of the Chueh-Conley-Smoller theory to establish uniform $L^{\infty}$ bounds via invariant regions. The convergence to a global weak solution in the vanishing regularization limit is finally proved using the method of compensated compactness and a singular analysis of the entropy equation near the vacuum.
@@ 1 Introduction
While inspired by Lu’s pioneering approach [4], we propose an alternative regularization technique that modifies only the pressure term, thereby strictly preserving the exact mass conservation law $\rho_{t}+(\rho u)_{x}=0$. To overcome the vacuum difficulty, we approximate (1.1) by the following perturbed system with a novel weighted pressure perturbation:
@@
This formulation ensures that the acoustic characteristics vanish exactly at $\rho=\delta$ (i.e., $\tilde{P}^{\prime}(\delta)=0$), creating a sequence of strictly hyperbolic systems. Our approach maintains the geometric structure of the invariant regions required for the $L^{\infty}$ estimates, but does so within a framework that respects the original continuity equation. The detailed motivation and construction of the perturbation function are provided in Appendix A, while a comparative analysis with Lu [4] is presented in Appendix B.
@@ 2.1 Properties of the Weighted Pressure
Defining the squared sound speed for the perturbed system as $\tilde{c}^{2}(\rho):=\partial_{\rho}\tilde{P}(\rho,\delta)$, we obtain from the definition (1.5):
$\tilde{c}^{2}(\rho)=P^{\prime}(\rho)-\frac{\delta^{2}P^{\prime}(\delta)}{\rho^{2}}.$ (2.2)
@@
1. Let $g(\rho)=\rho^{2}P^{\prime}(\rho)$. Compute $g^{\prime}(\rho)=\rho(2P^{\prime}+\rho P^{\prime\prime})>0$. Since $\rho>\delta$, monotonicity implies $g(\rho)>g(\delta)$, which gives $\rho^{2}\tilde{P}^{\prime}(\rho)=g(\rho)-g(\delta)>0$.
2. We compute the convexity term for the perturbed pressure. Using $\tilde{P}^{\prime}=\tilde{c}^{2}$:
@@
$=2P^{\prime}(\rho)-\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}+\rho P^{\prime\prime}(\rho)+\frac{2\delta^{2}P^{\prime}(\delta)}{\rho^{2}}$
@@
The singular terms cancel exactly. Since the original pressure satisfies $2P^{\prime}+\rho P^{\prime\prime}>0$, the perturbed pressure satisfies the exact same condition. ∎
@@ 2.2 Eigenstructure and Riemann Invariants
Let the vector of conservative variables be $\mathbf{U}=(\rho,m)^{T}$ with momentum $m=\rho u$. The flux vector is given by:
$\mathbf{F}(\mathbf{U})=\begin{pmatrix}m\\ \frac{m^{2}}{\rho}+\tilde{P}(\rho)\end{pmatrix}$. (2.3)
The system in quasi-linear form is $\mathbf{U}_{t}+A(\mathbf{U})\mathbf{U}_{x}=0$, where $A(\mathbf{U})=\nabla_{\mathbf{U}}\mathbf{F}$ is the Jacobian matrix. We compute the partial derivatives of the flux components $F_{1}=m$ and $F_{2}=m^{2}/\rho+\tilde{P}(\rho)$:
$\frac{\partial F_{1}}{\partial\rho}$ $=0,$ $\frac{\partial F_{1}}{\partial m}$ $=1,$
$\frac{\partial F_{2}}{\partial\rho}$ $=-\frac{m^{2}}{\rho^{2}}+\tilde{P}^{\prime}(\rho)=\tilde{c}^{2}-u^{2},$ $\frac{\partial F_{2}}{\partial m}$ $=\frac{2m}{\rho}=2u.$
Thus, the Jacobian matrix is:
$A(\mathbf{U})=\begin{pmatrix}0&1\\ \tilde{c}^{2}-u^{2}&2u\end{pmatrix}$. (2.4)
The eigenvalues $\lambda$ are determined by the characteristic equation $\det(A(\mathbf{U})-\lambda I)=0$:
$\det\begin{pmatrix}-\lambda&1\\ \tilde{c}^{2}-u^{2}&2u-\lambda\end{pmatrix}=-\lambda(2u-\lambda)-(\tilde{c}^{2}-u^{2})=\lambda^{2}-2u\lambda+(u^{2}-\tilde{c}^{2})=0.$ (2.5)
The eigenvalues are:
@@
The Riemann invariants $z,w$ are quantities constant along integral curves of the eigenvectors. They satisfy $du=\pm\frac{\tilde{c}(\rho)}{\rho}d\rho$. We define:
$z(\rho,u)=u-\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds,\quad w(\rho,u)=u+\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds.$ (2.6)
The integral is well-defined near $\rho=\delta$ because $\tilde{c}(\delta)=0$ and $(\tilde{c}^{2})^{\prime}(\delta)>0$, implying $\tilde{c}(s)\sim\sqrt{s-\delta}$ near the boundary.
Remark 2.2 (Consistency Limit).
It is straightforward to verify that as $\delta\to 0$, the perturbed invariants converge to the classical ones:
$\lim_{\delta\to 0}\int_{\delta}^{\rho}\frac{\tilde{c}(s)}{s}\,ds=\int_{0}^{\rho}\frac{\sqrt{P^{\prime}(s)}}{s}\,ds.$
This ensures that the invariant regions we construct later will converge to bounded regions for the original Euler system.
Proposition 2.3 (Explicit Form for Polytropic Gas).
For $P(\rho)=\kappa\rho^{\gamma}$, the Riemann invariants are:
$z(\rho,u)=u-\sqrt{\kappa\gamma}\int_{\delta}^{\rho}\frac{\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}}{s^{2}}\,ds,\quad w(\rho,u)=u+\sqrt{\kappa\gamma}\int_{\delta}^{\rho}\frac{\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}}{s^{2}}\,ds.$ (2.7)
@@
Substituting $P^{\prime}(s)=\kappa\gamma s^{\gamma-1}$ into (2.2) yields $\tilde{c}(s)=\frac{\sqrt{\kappa\gamma}}{s}\sqrt{s^{\gamma+1}-\delta^{\gamma+1}}$. The result follows by integration. ∎
Remark 2.4 (Comparison with Flux Modification).
It is worth noting why the mass flux modification in [4] is unnecessary here. Lu modified the continuity equation to $\rho_{t}+(\rho u-2\delta u)_{x}=0$ to force the eigenvalues to behave as $\lambda\approx u\pm\frac{\rho-2\delta}{\rho}\sqrt{P^{\prime}}$, thereby creating a degenerate boundary at $\rho=2\delta$. In contrast, our approach targets the sound speed directly. By constructing $\tilde{P}$ such that $\tilde{P}^{\prime}(\delta)=0$, we achieve the degeneracy $\lambda\to u$ at $\rho=\delta$ intrinsically. Furthermore, Lemma 2.1 shows that this construction preserves the convexity structure required for entropy analysis, rendering the auxiliary mass perturbation mathematically superfluous.
@@ 3 Entropy Analysis
To apply the compensated compactness method, we require a strictly convex entropy pair $(\eta,q)$ that satisfies the entropy identity $\partial_{t}\eta(\mathbf{U})+\partial_{x}q(\mathbf{U})=0$ for smooth solutions.
@@ 3.1 Construction of the Mechanical Energy Entropy Pair
The mechanical energy serves as the canonical entropy for the isentropic system. We define the specific internal energy $e(\rho)$ associated with the perturbed pressure via the thermodynamic relation $de=-\tilde{P}d(1/\rho)=\frac{\tilde{P}}{\rho^{2}}d\rho$. Normalizing such that $e(\delta)=0$, we obtain:
@@
The mechanical energy entropy (per unit volume) is defined as $\eta^{*}=\frac{1}{2}\rho u^{2}+\rho e(\rho)$. In terms of the conservative variables $(\rho,m)$ with $m=\rho u$, this reads:
@@
$q^{*}(\rho,m)=\left(\eta^{*}+\tilde{P}(\rho)\right)u=\left(\frac{m^{2}}{2\rho}+\rho e(\rho)+\tilde{P}(\rho)\right)\frac{m}{\rho}.$ (3.3)
@@
We verify the compatibility condition by computing the gradients with respect to the conservative variables $\mathbf{U}=(\rho,m)$. First, from the definition of $e(\rho)$, we have $e^{\prime}(\rho)=\frac{\tilde{P}(\rho)}{\rho^{2}}$. It follows that:
@@
Substituting these into the entropy equation for smooth solutions:
$\partial_{t}\eta^{*}+\partial_{x}q^{*}$ $=\eta^{*}_{\rho}\rho_{t}+\eta^{*}_{m}m_{t}+\partial_{x}q^{*}$
$=\left(-\frac{1}{2}u^{2}+e+\frac{\tilde{P}}{\rho}\right)(-\partial_{x}m)+u\left(-\partial_{x}(\rho u^{2}+\tilde{P})\right)+\partial_{x}\left((\eta^{*}+\tilde{P})u\right).$
A direct algebraic expansion confirms that the terms sum to zero. Thus, $(\eta^{*},q^{*})$ constitutes a valid entropy-flux pair. ∎
@@
The entropy $\eta^{*}(\rho,m)$ is strictly convex with respect to the conservative variables for $\rho>\delta$.
@@
We compute the Hessian matrix $\nabla^{2}\eta^{*}$:
@@
The second derivatives are given by:
$\eta^{*}_{mm}=\frac{\partial}{\partial m}\left(\frac{m}{\rho}\right)=\frac{1}{\rho},\quad\eta^{*}_{\rho m}=\frac{\partial}{\partial\rho}\left(\frac{m}{\rho}\right)=-\frac{m}{\rho^{2}}=-\frac{u}{\rho}.$
For $\eta^{*}_{\rho\rho}$, we differentiate $\eta^{*}_{\rho}$ with respect to $\rho$:
$\eta^{*}_{\rho\rho}$ $=\frac{\partial}{\partial\rho}\left(-\frac{u^{2}}{2}+e(\rho)+\frac{\tilde{P}}{\rho}\right)\quad(\text{noting }u=m/\rho)$
$=\frac{\partial}{\partial\rho}\left(-\frac{m^{2}}{2\rho^{2}}\right)+e^{\prime}(\rho)+\frac{\tilde{P}^{\prime}\rho-\tilde{P}}{\rho^{2}}$
$=\frac{m^{2}}{\rho^{3}}+\frac{\tilde{P}}{\rho^{2}}+\frac{\tilde{P}^{\prime}}{\rho}-\frac{\tilde{P}}{\rho^{2}}$
@@
$=\frac{u^{2}}{\rho^{2}}+\frac{\tilde{P}^{\prime}(\rho)}{\rho^{2}}-\frac{u^{2}}{\rho^{2}}$
@@
Since $\rho>\delta$, we have $\rho>0$. By Lemma 2.1, $\tilde{P}^{\prime}(\rho)=\tilde{c}^{2}(\rho)>0$ for $\rho>\delta$. Furthermore, the trace is $\text{tr}(\nabla^{2}\eta^{*})=\frac{u^{2}+\tilde{P}^{\prime}}{\rho}+\frac{1}{\rho}>0$. Positive trace and positive determinant imply that the Hessian is positive definite. Hence, $\eta^{*}$ is strictly convex for $\rho>\delta$. ∎
@@
This is a direct consequence of the constructed acoustic degeneracy at the cut-off density: $\tilde{P}^{\prime}(\delta)=0$. Consequently, the determinant of the Hessian vanishes at the boundary:
@@
The entropy $\eta^{*}$ therefore exhibits a degenerate convexity (or loss of strict convexity) at the boundary $\rho=\delta$, which geometrically corresponds to the flattening of the entropy surface.